# Hamiltonian Engineering of collective XYZ spin models in an optical cavity: From one-axis twisting to two-axis counter twisting models

### v1

@@ Introduction
We can derive the equations of motion of the collective Bloch vector, from Eq. (3) (see Method), which simplify to a nonlinear torque equation $d\mathbf{J}/dt=\mathbf{B}(\mathbf{J})\times\mathbf{J}\equiv\mathbf{T}(\mathbf{J})$. One can identify the fixed points $\mathbf{J}_{\rm fix}$ as the points where $\mathbf{T}(\mathbf{J}_{\rm fix})=0$.
To understand the dynamics near the fixed points, it is useful to follow a standard stability analysis by diagonalizing the Jacobian matrix $M(\mathbf{J})=\partial\mathbf{T}/\partial\mathbf{J}|_{\mathbf{J}=\mathbf{J}_{\rm fix}}$. The local motion near these fixed points is illustrated in Fig. 1(D) (right). We use red circles for stable points with purely imaginary eigenvalues. The Bloch vector evolves on stable closed orbits indicated by the blue circular traces. The red squares are used to denote unstable saddle points with real eigenvalues with opposite signs. The unstable saddle points are labelled by the red squares. The eigenvalues of the Jacobian matrix at the saddle points are real but with opposite signs. In Fig. 1(D) (bottom right), the dynamics show exponential divergence from the origin (indicated by outward blue arrows) along $\hat{x}+\hat{z}$, corresponding to the positive eigenvalue PoggiDeutsch2023PRXQuantum . The negative eigenvalues are indicated by the convergence towards the origin (inward blue arrows) along $\hat{x}-\hat{z}$.
Dynamics on the Bloch sphere
In the experiment, we probe the local dynamics induced by the above Hamiltonian with various values of $\chi_{e}$ and $\chi_{p}$. To do this, we vary the phase and duration of the Bragg pulse to prepare an initial pseudo-spin coherent state $\mathbf{J}_{i}$. Before the atomic wave packets separate, we then apply the interaction for a short time $\Delta t$ satisfying $\chi N\Delta t\ll 1$ and measure the change in azimuthal angle $d\phi$ and polar angle $d\theta$ to obtain the final Bloch vector $\mathbf{J}_{f}$ after the interaction. This is achieved by repeating the experiment and applying additional appropriate rotations before measuring populations in the two momentum states (see Methods). The local flow vector is then determined by the torque $\mathbf{T}(\mathbf{J}_{i})\approx\Delta\mathbf{J}/\Delta t=\left(\mathbf{J}_{f}-\mathbf{J}_{i}\right)/\Delta t$.
@@
We show the measured flow vector $\mathbf{T}(\mathbf{J}_{i})$ in Fig. 2(B) with each row aligned to the corresponding example presented in Fig. 2(A). In each case, the flow vector start at $\mathbf{J}_{i}$ and end at $\mathbf{J}_{f}$. The left and right panels are polar plots (radial coordinate linear in polar angle) of the dynamics on the south/north hemisphere looking from the north poles of the Bloch sphere. The middle panels are equirectangular projections to show dynamics near the equator. From these vector maps, we can make qualitative comparisons based on the geometry of the flow. For both theoretical and experimental results, the stable fixed points and unstable saddle points labeled as numbered red circles and squares respectively PoggiDeutsch2023PRXQuantum ; Witkowska2015TACTdynamics .
@@
In the first row of Fig. 2, we consider the simplest case with $\chi_{p}=0$. This is achieved by turning off one of the dressing lasers ($\left|\alpha_{2}/\alpha_{1}\right|=0$), leading to the Hamiltonian $\hat{H}=\chi_{e}\left(\hat{\bf J}\cdot\hat{\bf J}-\hat{J}_{z}^{2}\right)$. This Hamiltonian, referred to as the OAT Hamiltonian KitagawaUedaOAT , maintains U(1) symmetry, thereby conserving ${J}_{z}$. At the mean-field level, it features a constant effective magnetic field along the $\hat{z}$ direction, which results in the rotation of the collective Bloch vector about the $\hat{z}$-axis at a uniform constant angular frequency,$-2\chi_{e}{J}_{z}$. As expected, we observe two stable fixed points (red 1 and 2 circles), and a reversal of the circulation across the equator (i.e. $J_{z}=0$) where there are no dynamics (red line). This $J_{z}$ dependent circulation leads to shearing of the quantum noise in the orientation of a Bloch vector on the Bloch sphere, a semiclassical explanation for how OAT dynamics generate spin-squeezed states KitagawaUedaOAT ; Thompson2022SAI . We note that while the term $\hat{\bf J}\cdot\hat{\bf J}$ is trivial for our current observations at short times, at longer times, when inhomogeneities in our system manifest, it can lead to important dynamical effects as shown in Ref. Thompson2023MomentumExchange .
@@
Next, we consider the last row in Fig. 2 with $\chi_{p}=\chi_{e}$. This is achieved by using equal dressing laser amplitudes $\left|\alpha_{2}/\alpha_{1}\right|=1$. Here, the Hamiltonian is $\hat{H}=2\chi_{e}\hat{J}_{x}^{2}$, leading to OAT dynamics along the $\hat{x}$ direction. The corresponding dynamics are induced at the mean-field level by a magnetic field along $\hat{x}$ that preserves ${J}_{x}$ and induces a rotation about $\hat{x}$ with constant angular frequency $4\chi_{e}{J}_{x}$. It is noteworthy that the interaction strength here is twice that of the $\hat{J}^{2}_{z}$ case, attributed to the use of two dressing laser tones. The data qualitatively shows two stable fixed points along $\hat{x}$ labeled by red circles 1 and 2, and a reversal of the sign of circulation across ${J}_{x}$ highlighted by the red lines.
@@
We experimentally explore the flow lines. In the left and right panels, the observed stable fixed points are labeled with red circles numbered 1 to 4. The circulations of the flow lines are opposite for stable fixed points on opposite sides of the Bloch sphere. In the middle panel, we observe two unstable points labeled with red squared 1 and 2 in Fig. 2(C). The two unstable fixed points have flow lines that either diverge from them (blue arrows outward) or converge towards them (blue arrows inward), with the two flows orthogonal to each other (see the discussion below). These findings constitute the first direct observation of genuine TACT dynamics. We note that for a particular set of parameters of the LMG Hamiltonian, for example when $\hat{H}=\chi\hat{J}_{z}^{2}+\delta\hat{J}_{y}$ with $\delta\sim\chi N$, and when the Bloch vector initially points along the $\hat{y}$ direction PoggiDeutsch2023PRXQuantum ; Witkowska2015TACTdynamics , the flow lines can resemble the ones close to a saddle point of the TACT model (see Methods). However, the instability is restricted to this single point, in contrast to the full TACT which features two independent unstable points.
@@
Figure 3: Dynamics near saddle points. (A) Measured local flow vector map around the saddle point of the TACT dynamics. (B) Blue and red points are parallel projections of the local flow vector onto the $\hat{n}_{\pm}$ axis $\Delta\mathbf{J}\cdot\hat{n}_{\pm}/\left(N/2\right)$ as a function of initial displacement $d\mathbf{J}_{i}\cdot\hat{n}_{\pm}/\left(N/2\right)$ from the saddle point along the blue and red lines in (A). Blue and red solid lines are simulation results. The linear dependence shows exponentially growing / shrinking dynamics along the squeezing / anti-squeezing axes. (C) Green and orange points are the perpendicular projection of the local flow vector as a function of initial angular displacement from the saddle point along the green and orange lines in (A) ($\hat{n}=\hat{x}$ and $\hat{z}$) onto $\hat{n}\times\hat{y}=\hat{z}$ and $-\hat{x}$ axis ($\Delta\mathbf{J}_{z}$ and $-\Delta\mathbf{J}_{x}$). Orange and green solid lines are simulation results.
@@
For a quantitative comparison, we examine the dynamics near the saddle point $\mathbf{J}_{\rm sad}$ along the $\hat{y}$ axis on the Bloch sphere. In Fig. 3(A), we map the displacement $\mathbf{T}\left(\mathbf{J}_{i}\right)$ as a function of the initial Bloch vector orientation $\mathbf{J}_{i}=\mathbf{J}_{\rm sad}+d\mathbf{J}_{i}$. We scan the initial Bloch vector angles $\theta_{i}$ and $\phi_{i}$ over a range ±$\pm\pi/12$ centered about $\pi/2$ (i.e. about $\hat{y}$) with discrete points sampled using a detailed $11\times 1$ 11 grid. The mean-field equations of motion for the two orthogonal directions $\hat{n}_{\pm}=\left(\hat{x}\pm\hat{z}\right)/\sqrt{2}$ are:
@@
For the small range of angles around the $\hat{y}$ axis sampled in these measurements, one can assume ${J}_{y}\approx N/2$, and therefore find out that the displacement’s time derivative increases linearly with the displacement itself, indicating dynamics that change exponentially over time.
@@
The unstable dynamics explain why an initial circular distribution centered around the saddle point shears as a function of time by squeezing (anti-squeezing) along the $\hat{n}_{-}$ ($\hat{n}_{+}$) direction at a rate exponentially faster than the linear growth seen in OAT. This combined with the global (over the full Bloch sphere) dynamical behavior allows the TACT to directly approach the fundamental Heisenberg limit on phase estimation KitagawaUedaOAT ; PoggiDeutsch2023PRXQuantum .
@@
Two-axis counter-twisting with unstable points at north and south poles
The original TACT Hamiltonian, as proposed by Kitagawa and Ueda KitagawaUedaOAT , is defined as
@@
with the theoretical flow lines shown in Fig. 4(A). This Hamiltonian bears resemblance to the previously discussed TACT, with an additional $\pi/2$ rotation around the $y$-axis. It is characterized by unstable points at the north and south poles of the Bloch sphere, along with four stable points located on the equator at ±$\pm\hat{x}$ and ±$\pm\hat{y}$. In our setup, it seems to be impossible to choose the parameters canceling $\chi_{e}$ in Eq. (2) at first glance. However, this can be achieved by placing the two detuned dressing lasers tones on the same side of the cavity resonance as shown in Fig. 4(B) (top). In this configuration, the previously ignored lower modulation sideband of the red dressing laser becomes significant. This sideband introduces an exchange interaction with the opposite sign of the other generated exchange interactions. By carefully selecting the detuning and the amplitude ratio of the two dressing laser tones, we can achieve a configuration where $\chi_{e}=0$, effectively only leaving the $\chi_{p}$ term (see Methods.) $\hat{H}^{\prime}$ offers the advantage of possibly being less sensitive to the presence of superradiance or collective decay for future squeezing generation in our system.
Figure 4: Two-axis counter-twisting with unstable points at north/south poles. (A) Dynamics shown as flows on the Bloch sphere with three characteristic circles (purple: $\theta_{i}=0.9\pi$, black: $\theta_{i}=0.5\pi$ and green: $\theta_{i}=0.1\pi$) highlighted. (B) Frequency diagram with both dressing lasers positioned on the same side relative to the cavity resonance (top). The equirectangular projection of the resulting dynamics (bottom). (C), (D) and (E) The distribution of the states after the interaction with the prepared initial state $\theta_{i}=0.1\pi$ (green circle), $0.9\pi$ (purple circle) and $0.5\pi$ (black circle). The deviations from the initial state distribution (gray dashed circles) are shown by the colored arrows.
@@ Conclusion
Here we have demonstrated the flexibility of our optical cavity simulator to engineer tunable XYZ Hamiltonians using two selected momentum states, without the need for Floquet engineering which in some cases might be challenging to allow access beyond mean-field dynamics in large many-body systems due to amplitude and phase noise on the Floquet control fields ThompsonPhaseNoise2012 . The pair-raising/lowering processes are verified to be present for the first time by the resonant spectroscopic signal when the detuning $\delta$ is scanned as well as the direct cancellation of the exchange interactions that yielded dynamics of the Hamiltonian $\hat{H}^{\prime}=\chi\left(\hat{J}_{+}^{2}+\hat{J}_{-}^{2}\right)$. By combining the correct relative balance of exchange and pair-raising/lowering contributions, we observed TACT dynamics for the first time.
While so far our focus has been on the characterization of the Hamiltonian by probing the short-time dynamics, extension to longer times can open untapped opportunities for quantum simulation and metrology. For example, by using an implementation where two dressing lasers are used to generate one-axis twisting dynamics, one should be able to balance the collective dissipation generated by the absorbed and emitted photons facilitating the emergence of beyond mean-field effects and entanglement generation.

### v2

@@ III Mean-field dynamics
We can derive the equations of motion of the collective Bloch vector, from Eq. (LABEL:eq:heff) (see Method), which simplify to a nonlinear torque equation $d\mathbf{J}/dt=\mathbf{B}(\mathbf{J})\times\mathbf{J}\equiv\mathbf{T}(\mathbf{J})$. One can identify the fixed points $\mathbf{J}_{\rm fix}$ as the points where $\mathbf{T}(\mathbf{J}_{\rm fix})=0$.
To understand the dynamics near the fixed points, it is useful to follow a standard stability analysis by diagonalizing the Jacobian matrix $M(\mathbf{J})=\partial\mathbf{T}/\partial\mathbf{J}|_{\mathbf{J}=\mathbf{J}_{\rm fix}}$. The local motion near these fixed points is illustrated in Fig. 1(D) (right). We use red circles for stable points with purely imaginary eigenvalues. The Bloch vector evolves on stable closed orbits indicated by the blue circular traces. The red squares are used to denote unstable saddle points with real eigenvalues with opposite signs. The unstable saddle points are labelled by the red squares. The eigenvalues of the Jacobian matrix at the saddle points are real but with opposite signs. In Fig. 1(D) (bottom right), the dynamics show exponential divergence from the origin (indicated by outward blue arrows) along $\hat{x}+\hat{z}$, corresponding to the positive eigenvalue [47]. The negative eigenvalues are indicated by the convergence towards the origin (inward blue arrows) along $\hat{x}-\hat{z}$.
IV Dynamics on the Bloch sphere
In the experiment, we probe the local dynamics induced by the above Hamiltonian with various values of $\chi_{e}$ and $\chi_{p}$. To do this, we vary the phase and duration of the Bragg pulse to prepare an initial pseudo-spin coherent state $\mathbf{J}_{i}$. Before the atomic wave packets separate, we then apply the interaction for a short time $\Delta t=50\mu$s satisfying $\chi N\Delta t\ll 1$ and measure the change in azimuthal angle $d\phi$ and polar angle $d\theta$ to obtain the final Bloch vector $\mathbf{J}_{f}$ after the interaction. This is achieved by repeating the experiment and applying additional appropriate rotations before measuring populations in the two momentum states (see supplement). The local flow vector is then determined by the torque $\mathbf{T}(\mathbf{J}_{i})\approx\Delta\mathbf{J}/\Delta t=\left(\mathbf{J}_{f}-\mathbf{J}_{i}\right)/\Delta t$.
@@ IV Dynamics on the Bloch sphere
We show the measured flow vector $\mathbf{T}(\mathbf{J}_{i})$ in Fig. 2(B) with each row aligned to the corresponding example presented in Fig. 2(A). In each case, the flow vector start at $\mathbf{J}_{i}$ and end at $\mathbf{J}_{f}$. The left and right panels are polar plots (radial coordinate linear in polar angle) of the dynamics on the south/north hemisphere looking from the north poles of the Bloch sphere. The middle panels are equirectangular projections to show dynamics near the equator. From these vector maps, we can make qualitative comparisons based on the geometry of the flow. For both theoretical and experimental results, the stable fixed points and unstable saddle points labeled as numbered red circles and squares respectively [47, 48].
IV.1 OAT: ${\hat{J}_{z}^{2}}$ interaction
@@
In the first or top row of Fig. 2 , we consider the simplest case with $\chi_{p}=0$. This is achieved by turning off one of the dressing lasers ($\left|\alpha_{2}/\alpha_{1}\right|=0$), leading to the Hamiltonian $\hat{H}=\chi_{e}\left(\hat{\bf J}\cdot\hat{\bf J}-\hat{J}_{z}^{2}\right)$. This Hamiltonian, referred to as the OAT Hamiltonian [19], maintains U(1) symmetry, thereby conserving ${J}_{z}$. At the mean-field level, it features a constant effective magnetic field along the $\hat{z}$ direction, which results in the rotation of the collective Bloch vector about the $\hat{z}$-axis at a uniform constant angular frequency,$-2\chi_{e}{J}_{z}$. As expected, we observe in Fig. 2(B) two stable fixed points (red 1 and 2 circles), and a reversal of the circulation across the equator (i.e. $J_{z}=0$) where there are no dynamics (red line). This $J_{z}$ dependent circulation leads to shearing of the quantum noise in the orientation of a Bloch vector on the Bloch sphere, a semiclassical explanation for how OAT dynamics generate spin-squeezed states [19, 35]. We note that while the term $\hat{\bf J}\cdot\hat{\bf J}$ is trivial for our current observations at short times, at longer times, when inhomogeneities in our system manifest, it can lead to important dynamical effects as shown in Ref. [46].
IV.2 OAT: ${\hat{J}_{x}^{2}}$ interaction
@@
Next, we consider the bottom row in Fig. 2 with $\chi_{p}=\chi_{e}$. This is achieved by using equal dressing laser amplitudes $\left|\alpha_{2}/\alpha_{1}\right|=1$. Here, the Hamiltonian is $\hat{H}=2\chi_{e}\hat{J}_{x}^{2}$, leading to OAT dynamics along the $\hat{x}$ direction. The corresponding dynamics are induced at the mean-field level by a magnetic field along $\hat{x}$ that preserves ${J}_{x}$ and induces a rotation about $\hat{x}$ with constant angular frequency $4\chi_{e}{J}_{x}$. It is noteworthy that the interaction strength here is twice that of the $\hat{J}^{2}_{z}$ case, attributed to the use of two dressing laser tones. The data in Fig. 2(D) qualitatively shows two stable fixed points along $\hat{x}$ labeled by red circles 1 and 2, and a reversal of the sign of circulation across ${J}_{x}$ highlighted by the red lines.
IV.3 TACT: $\hat{J}_{x}^{2}-\hat{J}_{z}^{2}$ interaction
@@
In the left and right panels of Fig. 2(C), the observed stable fixed points are labeled with red circles numbered 1 to 4. The circulations of the flow lines are opposite for stable fixed points on opposite sides of the Bloch sphere. In the middle panel, we observe two unstable points labeled with red squared 1 and 2 in Fig. 2(C). The two unstable fixed points have flow lines that either diverge from them (blue arrows outward) or converge towards them (blue arrows inward), with the two flows orthogonal to each other (see the discussion below). These findings constitute the first direct observation of genuine TACT dynamics. We note that for a particular set of parameters of the LMG Hamiltonian, for example when $\hat{H}=\chi\hat{J}_{z}^{2}+\delta\hat{J}_{y}$ with $\delta\sim\chi N$, and when the Bloch vector initially points along the $\hat{y}$ direction [47, 48], the flow lines can resemble the ones close to a saddle point of the TACT model (see supplement). However, the instability is restricted to this single point, in contrast to the full TACT which features two independent unstable points.
The observed dynamics reflect the different symmetries of the TACT and OAT Hamiltonians. Both of the observed OAT dynamics are symmetric under $\pi$ rotation around $\hat{y}$ and have continuous rotational symmetry about the twisting axis. In contrast, the flow lines for the TACT dynamics, shown in the middle row, does not have such symmetries. The TACT dynamics nevertheless, is symmetric under a $\pi/2$ rotation along $\hat{y}$ plus a time-reversal operation.
@@
Figure 3: Dynamics near saddle points. (A) Measured local flow vector map around the saddle point of the TACT dynamics. (B) Blue and red points are parallel projections of the local flow vector onto the $\hat{n}_{\pm}$ axis $\Delta\mathbf{J}\cdot\hat{n}_{\pm}/\left(N/2\right)$ as a function of initial displacement $d\mathbf{J}_{i}\cdot\hat{n}_{\pm}/\left(N/2\right)$ from the saddle point along the blue and red lines in (A). Blue and red solid lines are simulation results. The linear dependence confirms the predicted exponentially growing / shrinking dynamics along the squeezing / anti-squeezing axes. (C) Green and orange points are the perpendicular projection of the local flow vector as a function of initial angular displacement from the saddle point along the green and orange lines in (A) ($\hat{n}=\hat{x}$ and $\hat{z}$) onto $\hat{n}\times\hat{y}=\hat{z}$ and $-\hat{x}$ axis ($\Delta\mathbf{J}_{z}$ and $-\Delta\mathbf{J}_{x}$). Orange and green solid lines are simulation results.
@@
For a quantitative comparison, we examine the dynamics near the saddle point $\mathbf{J}_{\rm sad}$ along the $\hat{y}$ axis on the Bloch sphere. In Fig. 3(A), we map the displacement $\mathbf{T}\left(\mathbf{J}_{i}\right)$ as a function of the initial Bloch vector orientation $\mathbf{J}_{i}=\mathbf{J}_{\rm sad}+d\mathbf{J}_{i}$. We scan the initial Bloch vector angles $\theta_{i}$ and $\phi_{i}$ over a range ±$\pm\pi/12$ centered about $\pi/2$ (i.e. about $\hat{y}$) with discrete points sampled using a detailed $11\times 1$ 11 grid. In Fig. 3, we remove a global displacement along $\hat{z}$ that is common to all data points and which arises from collective or superradiant decay (see supplement.)
The mean-field equations of motion for the two orthogonal directions $\hat{n}_{\pm}=\left(\hat{x}\pm\hat{z}\right)/\sqrt{2}$ are:
@@
For the small range of angles around the $\hat{y}$ axis sampled in these measurements, one can assume ${J}_{y}\approx N/2$, and therefore find out that the displacement’s time derivative increases linearly with the displacement itself, confirming the predicted dynamics that change exponentially over time.
@@
The unstable dynamics explain why an initial circular distribution centered around the saddle point shears as a function of time by squeezing (anti-squeezing) along the $\hat{n}_{-}$ ($\hat{n}_{+}$) direction at a rate exponentially faster than the linear growth seen in OAT. This combined with the global (over the full Bloch sphere) dynamical behavior allows the TACT to directly approach the fundamental Heisenberg limit on phase estimation [19, 47].
@@
V Two-axis counter-twisting with unstable points at north and south poles
The original TACT Hamiltonian, as proposed by Kitagawa and Ueda [19], is defined as
@@ V Two-axis counter-twisting with unstable points at north and south poles
Figure 4: Two-axis counter-twisting with unstable points at north/south poles. (A) Flow lines on the Bloch sphere. Three characteristic closed trajectories are displayed (green: $\theta_{i}=0.1\pi$, purple: $\theta_{i}=0.9\pi$, and black: $\theta_{i}=0.5\pi$ ) (B) (top) Frequency diagram for the case in which both dressing lasers are positioned on the same side of the cavity resonance. (bottom) Equirectangular projection of the resulting dynamics (see text) (C), (D) and (E) Spin projections obtained after time evolution of an initial distribution of Bloch vectors with different polar angles (green circle: $\theta_{i}=0.1\pi$, purple circle: $\theta_{i}=0.9\pi$ and black circle: $0.5\pi$ ). The deviation from the initial state distribution (gray dashed circles) is indicated by the colored arrows.
with the theoretical flow lines shown in Fig. 4(A). This Hamiltonian bears resemblance to the previously discussed TACT, with an additional $\pi/2$ rotation around the $y$-axis. It is characterized by unstable points at the north and south poles of the Bloch sphere, along with four stable points located on the equator at ±$\pm\hat{x}$ and ±$\pm\hat{y}$. We realize this Hamiltonian by placing the two detuned dressing lasers tones on the same side of the cavity resonance with $\Delta_{\bar{c}}=2\pi\times 700~{}\mathrm{kHz}>\omega_{z}$, such that one needs to account for another modulation sideband with opposite detuning from the cavity resoannce as shown in Fig. 4(B) (top). In this configuration, the previously ignored lower modulation sideband of the red dressing laser becomes significant. This sideband introduces an exchange interaction with the opposite sign of the other generated exchange interactions. By carefully selecting the detuning and the amplitude ratio of the two dressing laser tones, we can achieve a configuration where $\chi_{e}=0$, effectively only leaving the $\chi_{p}$ term (see supplement.) $\hat{H}^{\prime}$ offers the advantage of possibly being less sensitive to the presence of superradiance or collective decay for future squeezing generation in our system.
@@ VI Conclusion
Here we have demonstrated the flexibility of our optical cavity simulator to engineer tunable XYZ Hamiltonians using two selected momentum states, without the need for Floquet engineering which in some cases might be challenging to allow access beyond mean-field dynamics in large many-body systems due to amplitude and phase noise on the Floquet control fields [49]. The pair-raising/lowering processes are verified to be present for the first time by the resonant spectroscopic signal when the detuning $\delta$ is scanned as well as the direct cancellation of the exchange interactions that yielded dynamics of the Hamiltonian $\hat{H}^{\prime}=\chi\left(\hat{J}_{+}^{2}+\hat{J}_{-}^{2}\right)$. By combining the correct relative balance of exchange and pair-raising/lowering contributions, we observed TACT dynamics for the first time.
In this work, we focus on the characterization of the Hamiltonian by probing the short-time dynamics at the mean-field level. In the future, it should be possible to go beyond mean field and create entangled states of interest for quantum simulation and metrology. Going beyond mean field will require reduced laser frequency noise on the dressing lasers that induce the interaction.